# Third law of repetitive electric Penrose process

### v1

@@ Abstract
Recently, Ruffini et al. [Phys. Rev. Lett. 134 (2025) 8, 081403] pointed out that the repetitive Penrose process cannot drain the entire extractable energy of a Kerr black hole. In this Letter, we further point out that the charge of a Reissner-Nordström black hole cannot drop down to exactly zero via the repetitive electric Penrose process that is terminated after a finite number of iterative steps, indicating a thermodynamical third-law analog for the repetitive electric Penrose process.
Introduction. In the original proposal of Penrose process Penrose and Floyd (1971), an incoming particle decays within the ergosphere of a Kerr black hole into a negative-energy particle that falls into the black-hole horizon and a positive-energy particle that escapes to infinity. Due to the conservation of energy, the energy of escaping particle is always greater than that of incoming particle, allowing us to extract energy from the rotating black hole. Soon after, the electric Penrose process (EPP) was proposed Denardo and Ruffini (1973) by replacing the Kerr background with a Reissner-Nordström (RN) geometry, and finding that the negative-energy orbit also exists in the dubbed generalized ergosphere, making it feasible to extract energy from a charged black hole through the electromagnetic version of the Penrose process Mai and Yang (2023).
@@
In this letter, we turn to the repetitive EPP in the RN black hole. We find a similar incapability of extracting all extractable energy, and the charge of the RN black hole can be very small eventually, but never vanishes solely from the EPP that absorbs a decay product particle with negative energy. The charge of an RN black hole can only possibly vanish either from the semiclassical/quantum process by Hawking evaporation Brown et al. (2024) or accidental neutralization by absorbing a positive-energy particle with exactly equal but opposite charge.
@@
Recall that in Ref. Zaslavskii (2024) the decay products are ejected along the same trajectory of the incident particle to achieve a maximum efficiency. To avoid the complication of Kerr-Newman geometry, we further simplify the configuration of the incident particle from infinity, either at rest or with an initial radial velocity Kokubu et al. (2021), so that both particles 1 and 2 maintain radial motion to maximize the energy return on investment (EROI) (defined later).
Besides the conservation equations (2), (3), and (4), the normalization of the proper four velocity gives rise to another three equations,
@@
for $i=0,1,2$, where $\hat{V}_{i}^{\pm}$ is understood as the effective potential of the radial motion Ruffini and Wheeler (1971). Now we have 6 equations and 11 variables (2 masses, 3 energies, 3 charges, 3 radial momenta). However, the maximum EROI condition can further eliminate the equation (4). Refs. Ruffini et al. (2025a, b) show that in the Penrose process, the maximum EROI corresponds to the decay when all three particles are at their own turning points Ruffini et al. (2025b) with vanishing kinetic energy, that is, the radial momenta of the three particles must vanish at the decay location. We will show shortly below that this correspondence still holds for EPP.
@@
$\Big(\frac{\mathrm{d}t}{\mathrm{d}\tau}\Big)_{1}=\frac{\hat{E}_{1}-\frac{\hat{Q}\hat{q}_{1}}{\hat{r}}}{\sqrt{f}}>0\,.$ (9)
Therefore, all three particles must be located at their turning points with vanishing kinetic energy at the decay location to maximize the efficiency, and hence eliminating the equation (4).
@@
$\hat{E}_{2}$ $=\frac{\hat{Q}\hat{q}_{2}}{\hat{r}_{d}}+\sqrt{1-\frac{2}{\hat{r}_{d}}+\frac{\hat{Q}^{2}}{\hat{r}_{d}^{2}}}\,.$ (14)
@@
First, the classical turning points of particles 0 and 2 must lie outside the peaks of their effective potentials, while the classical turning point of particle 1 must lie inside the peak of its effective potential. The corresponding limiting case is when the classical turning points of each particle lie exactly at the peak of each effective potential, $\hat{V}_{i}^{+}(\hat{r}_{d})=\hat{E}_{i}$, and $\mathrm{d}\hat{V}_{i}^{+}/\mathrm{d}\hat{r}|_{\hat{r}=\hat{r}_{d}}=0$. Solving above two equations simultaneously yields the minimal value of $\hat{Q}$,
$\hat{Q}^{2}_{\mathrm{min}}=\frac{(\hat{r}_{d}-1)^{2}}{\hat{E}_{i}^{2}}+\hat{r}_{d}(2-\hat{r}_{d})\,,\quad i=0,1,2\,.$ (17)
@@
Consequently, the overall EUE stands at only $22\%$, with nearly three-quarters of the reduced extractable energy being converted into the irreducible mass of the black hole (that is, the increase in entropy). This portion is thereby permanently locked up according to Hawking’s area law, imposing a fundamental thermodynamic cost on the energy extraction process.
@@
Figure 3: Variation of EROI and EUE of the final state with the dimensionless decay radius, the charge-to-mass ratio of particle 1, and the reduced energy of particle 0. Here we keep the mass ratio $\eta=0.78345$ but choose a smaller mass of the incident particle $\mu_{0}=M_{0}/1000$ to eliminate the “oscillation” of the curves as much as possible. The basic choice of three variables is $\hat{r}_{d}=2.4$, $\hat{q}_{1}=-5$, and $\hat{E}_{0}=1.1$. Each time one of the three variables is changed while keeping the other two fixed. It is intuitive to see that $\xi_{n_{\mathrm{end}}}$ can easily be much greater than $1$, but $\Xi_{n_{\mathrm{end}}}$ is difficult to exceed $0.5$.
@@
It is straightforward to find that both EROI and EUE decrease significantly as the decay location moves further away from the black hole horizon, while both of them increase as the charge-to-mass ratio of particle 1 increases. Regarding the velocity of the incident particle, since the energy extracted by a single process is nearly “fixed” by $\hat{q}_{1}$, EROI decreases as $\hat{E}_{0}$ becomes bigger. Interestingly, EUE stops growing and becomes constant when $\hat{E}_{0}$ exceeds $1.4$. This is because the stopping condition switches from the dashed curve in Fig. 1 to the solid curve when $\hat{E}_{0}>1.4$, which means that the final state of the black hole is fixed and will no longer change with $\hat{E}_{0}$. One thing to note is that we do not show the case $|\hat{q}_{1}|<1$ here because it may not be an extreme black hole initially, making it difficult to compare it with other cases directly.
Conclusions and discussions. We study the detailed behavior of the repetitive EPP under the condition of maximizing the EROI. We first find that the parameter choices (the dimensionless decay position and the reduced black hole charge) are constrained by both the incident particle velocity and the charge-to-mass ratio of the product particle that falls into the black hole. Given the substantially high charge-to-mass ratio of microscopic particles and the abundance of high-energy cosmic particles in the Universe, the charge of the RN black hole can be efficiently neutralized to a negligibly small value through repetitive EPP. Nevertheless, it is noteworthy that the charge cannot be fully discharged to zero via this classical process.
@@
Acknowledgments. This work is supported by the National Key Research and Development Program of China Grant No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12588101, No. 12235019, and No. 12447101, and the China Manned Space Program Grant No. CMS-CSST-2025-A01.
@@ I.1 Improvement to previous work
The constraint obtained in Ruffini et al. Ruffini et al. (2025a, b) can be further improved by including the second stopping condition (it might be more accurate to call it the starting condition here) $\hat{E}_{1}<0$, rendering $\hat{a}<\sqrt{\frac{2\hat{p}_{\phi 1}^{2}-\hat{p}_{\phi 1}^{2}\hat{r}^{2}+2\hat{r}^{2}-\hat{r}^{3}}{\hat{r}}}$ on the equatorial plane, which has further shrunk the allowed region in the parameter space as shown in Fig. 4.

### v2

@@ Abstract
Recently, Ruffini et al. [Phys. Rev. Lett. 134 (2025) 8, 081403] pointed out that the repetitive Penrose process cannot drain the entire extractable energy of a Kerr black hole. In this Letter, we alternatively point out the charge of a Reissner-Nordström black hole cannot drop down to exactly zero via the repetitive electric Penrose process that is terminated after a finite number of iterative steps, indicating a new thermodynamical third-law analog for the repetitive electric Penrose process.
Introduction. In the original proposal of Penrose process Penrose and Floyd (1971), an incoming particle decays within the ergosphere of a Kerr black hole into a negative-energy particle that falls into the black-hole horizon and a positive-energy particle that escapes to infinity. Due to the conservation of energy, the energy of escaping particle is always greater than that of incoming particle, allowing us to extract energy from the rotating black hole. Soon after, the electric Penrose process (EPP) was proposed Denardo and Ruffini (1973) by replacing the Kerr background with a Reissner-Nordström (RN) geometry, and finding that the negative-energy orbit also exists in the dubbed generalized ergosphere, making it feasible to extract energy from a charged black hole through the electromagnetic version of the Penrose process Mai and Yang (2023); Tursunov et al. (2021).
@@
In this letter, we turn to the repetitive EPP in the RN black hole. We find a similar incapability of extracting all extractable energy, and the charge of the RN black hole can be very small eventually, but never vanishes solely from the EPP that absorbs a decay product particle with negative energy. The charge of an RN black hole can only possibly vanish either from the semiclassical/quantum process by Hawking evaporation Brown et al. (2024) or accidental neutralization by absorbing positive-energy particles with exactly equal but opposite charge.
@@
Recall that in Ref. Zaslavskii (2024) the decay products are ejected along the same trajectory of the incident particle to achieve a maximum efficiency. Since we are dealing with the RN metric, we further simplify the configuration of the incident particle from infinity, either at rest or with an initial radial velocity Kokubu et al. (2021), so that both particles 1 and 2 maintain radial motion to maximize the energy return on investment (EROI) (defined later).
Besides the conservation Eqs. (2), (3), and (4), the normalization of the proper four velocity gives rise to another three equations,
@@
for $i=0,1,2$, where $\hat{V}_{i}^{\pm}$ is understood as the effective potential of the radial motion Ruffini and Wheeler (1971), and $\hat{r}\equiv r/M$ is the dimensionless radius. Now we have 6 equations and 11 variables (2 masses, 3 energies, 3 charges, 3 radial momenta). However, the maximum EROI condition can further eliminate the Eq. (4). Refs. Ruffini et al. (2025a, b) show that in the Penrose process, the maximum EROI corresponds to the decay when all three particles are at their own turning points Ruffini et al. (2025b) with vanishing kinetic energy, that is, the radial momenta of the three particles must vanish at the decay location. We will show shortly below that this correspondence still holds for EPP.
@@
$\Big(\frac{\mathrm{d}t}{\mathrm{d}\tau}\Big)_{1}=\frac{\hat{E}_{1}-\hat{Q}\hat{q}_{1}/\hat{r}}{\sqrt{f}}>0\,$, (9)
where $\tau$ is the proper time along the geodesic of particle 1. Therefore, all three particles must be located at their turning points with vanishing kinetic energy at the decay location to maximize the efficiency, and hence eliminating Eq. (4).
@@
$\hat{E}_{2}$ $=\frac{\hat{Q}\hat{q}_{2}}{\hat{r}_{d}}+\sqrt{1-\frac{2}{\hat{r}_{d}}+\frac{\hat{Q}^{2}}{\hat{r}_{d}^{2}}}\,$, (14)
where $\hat{r}_{d}$ denotes the dimensionless decay radius.
@@
$\hat{Q}^{2}_{i,\mathrm{min}}=\frac{(\hat{r}_{d}-1)^{2}}{\hat{E}_{i}^{2}}+\hat{r}_{d}(2-\hat{r}_{d})\,,\quad i=0,2\,.$ (17)
Since $E_{2}>E_{0}$ implies $\hat{E}_{2}>\hat{E}_{0}$, the constraint from particle 2 is weaker. Thus, the lower limit on $\hat{Q}$ is actually determined by particle 0,
@@
Consequently, the overall EUE stands at only $22\%$, with nearly three-quarters of the reduced extractable energy being converted into the irreducible mass of the black hole (that is, the increase in entropy). This portion is thereby permanently locked up according to Hawking’s area law, imposing a fundamental thermodynamic cost on the energy extraction process (see also Ref. Bravetti et al. (2016)).
@@
Figure 3: Variation of EROI and EUE of the final state with the dimensionless decay radius, the charge-to-mass ratio of particle 1, and the reduced energy of particle 0. Here we keep the mass ratio $\eta=0.78345$ but choose a smaller mass of the incident particle $\mu_{0}=M_{0}/1000$ to eliminate the “oscillation” of the curves as much as possible. The basic choice of three variables is $\hat{r}_{d}=2.4$, $\hat{q}_{1}=-5$, and $\hat{E}_{0}=1.1$. Each time one of the three variables is changed while keeping the other two fixed. It is intuitive to see that $\xi_{n_{\mathrm{end}}}$ can easily be much greater than 11, but $\Xi_{n_{\mathrm{end}}}$ is difficult to exceed $0.5$.
@@
It is straightforward to find that both EROI and EUE decrease significantly as the decay location moves further away from the black hole horizon, while both of them increase as the charge-to-mass ratio of particle 1 increases. Regarding the velocity of the incident particle, since the energy extracted by a single process is nearly “fixed” by $\hat{q}_{1}$, EROI decreases as $\hat{E}_{0}$ becomes bigger. Interestingly, EUE stops growing and becomes constant when $\hat{E}_{0}$ exceeds $1.4$. This is because the stopping condition switches from the dashed curve in Fig. 1 to the solid curve when $\hat{E}_{0}>1.4$, that is, the final state of the black hole is fixed and will no longer change with $\hat{E}_{0}$. One thing to note is that we do not show the case $|\hat{q}_{1}|<1$ here because it may not be an extreme black hole initially, making it difficult to compare it with other cases directly.
Conclusions and discussions. We look into the repetitive EPP under the condition of maximizing the EROI. We first find the parameter choices (the dimensionless decay position and reduced BH charge) are constrained by both the incident particle velocity and the charge-to-mass ratio of the product particle that falls into the black hole. Given the substantially high charge-to-mass ratio of microscopic particles and the abundance of high-energy cosmic particles in the Universe, the charge of the RN black hole can be efficiently neutralized to a negligibly small value through repetitive EPP. Nevertheless, it is noteworthy that the charge cannot be fully discharged to zero via this classical process, revealing another fundamental limit in a classic BH process.
@@
Acknowledgments. This work is supported by the National Key Research and Development Program of China Grants No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12547110, No.12588101, No. 12235019, and No. 12447101, and the China Manned Space Program Grant No. CMS-CSST-2025-A01.
@@ I.1 Improvement to previous work
The constraint obtained in Ruffini et al. Ruffini et al. (2025a, b) can be further improved by including the second stopping condition (it might be more accurate to call it the starting condition here) $\hat{E}_{1}<0$, rendering $\hat{a}<\sqrt{\frac{2\hat{p}_{\phi 1}^{2}-\hat{p}_{\phi 1}^{2}\hat{r}^{2}+2\hat{r}^{2}-\hat{r}^{3}}{\hat{r}}}$ on the equatorial plane, which has further shrunk the allowed region in the parameter space as shown in Fig. 4. The spin parameter $\hat{a}$ is defined in terms of the angular momentum $L$ as $\hat{a}\equiv a/M=L/M^{2}$, while the dimensionless angular momentum of particle 1 is given by $\hat{p}_{\phi 1}\equiv p_{\phi 1}/(\mu_{1}M)$.